OpenStudy (anonymous):

Evaluate the integral by using multiple substitutions ∫5(3x^2-7)sin^5(x^3-7x) cos(x^3-7x) dx

OpenStudy (anonymous):

\[\int\limits 5(3x^2-7)\sin^5(x^3-7x)\cos(x^3-7x)dx\]

OpenStudy (anonymous):

u= x^3−7x du = 3x^2-7 dx

OpenStudy (anonymous):

well \[u=x^3-7x;du=3x^2-7\] this leaves \[\int5sin^5(u)cos(u)du\]

OpenStudy (anonymous):

v =sin (u) dv = cos(u) du

OpenStudy (anonymous):

so the answer would be \[5sin^6(x^3-7x)\]

OpenStudy (anonymous):

Makes sense... THANK YOU BOTH!! ^_^

OpenStudy (anonymous):

T.T i thought i could give both of u medals T.T